OpenStudy (anonymous):

What is the solution of the system? use substitution. y=x+6y=2x a. (6,12) b. (-12,-6) c. (-6,-12) d. (2,4) IM VERY CONFUSED so please explain to me how to get to the answer.

OpenStudy (anonymous):

You should have two equations, did you copy them wrong?

OpenStudy (anonymous):

thats what it says.. here is the screen shot of it

OpenStudy (anonymous):

I've tried but sorry I cant help =( Good luck though, but I appear to be going in circles

OpenStudy (anonymous):

well thanks i guess the test question is wrong.

OpenStudy (anonymous):

i called the school is supposed to be y=x+6 y=2x can you help me now?

OpenStudy (anonymous):

y = x + 6 y = 2x And you've to work out what x and y are Okey looks like similtaeneous equasions If I write; y - x = 6 (HINT: -2y + 2x = -12) y - 2x = 0 Will you be abe to get that one yourself?

OpenStudy (anonymous):

For the substitution method, you solve one equation for one variable, then you substitute that into the other equation. if y=x+6 (first equation) then in the second equation y=2x you substitute that (x+6)=2x Then solve it 2x-x-6=0 x-6=0 x=6 Now, since y=x+6, y=12 Cool?